OpenStudy (anonymous):

(x+3/3)+(x-2/4)=1 answer is -6/7

OpenStudy (anonymous):

(x+3/3) +(x-2/4)=1 The answer I came up with is 1/4. Are you sure the answer is -6/7

OpenStudy (anonymous):

thats what I was told the answer was

OpenStudy (anonymous):

(x+3)/3 +(x-2)/4 =1  x = 6/7